# Facial Expression Analysis and Its Potentials in IoT Systems: A Contemporary Survey

### v1

@@ 5.2. MiE Recognition
Since the seminar work (Pfister et al., 2011a), MiE recognition has seen unprecedented development that ranges from shallow machine learning with handcrafted descriptors to deep learning. Hereinafter of this part, the MiE recognition will be thoroughly reviewed.
@@ 5.2.1. Shallow machine learning methods
The success of shallow machine learning [^1] The word “shallow” is used to differentiate from the deep learning methods. demands for the well-designed handcrafted descriptors (e.g., LBPs and grayscale images) to extract local and global facial information. As a key component of MiE recognition, the development of descriptors has attracted several research endeavors, such as, LBP with six intersection points (Wang et al., 2015b) and LBP with three mean orthogonal planes (Wang et al., 2015a). More specifically, Wang et al. (Wang et al., 2015b) proposed the LBP with six intersection points to reduce the redundancy inherent in LBP-TOP. Later, Wang et al. (Wang et al., 2015a) proposed the LBP with three mean orthogonal planes to alleviate the computational complexity issue of LBP-TOP. Compared with the LBP-TOP descriptor, the proposed LBP with three mean orthogonal planes offers faster processing with a competitive performance (Wang et al., 2015a). Huang et al. (Huang et al., 2016) developed the spatio-temporal completed local quantization pattern descriptor that incorporates additional information on sign, magnitude, and orientation components to enhance MiE recognition. To address the limitation of LBP-TOP on capturing muscle movement in the oblique direction, Wei et al. (Wei et al., 2022a) proposed a descriptor named as LBP with five intersecting planes (LBP-FIP). The LBP-FIP descriptor enhances the LBP-TOP with an innovative feature (i.e., eight vertices LBP) in order to enrich the feature information by capturing detailed MiE dynamics in multiple directions. Therefore, the LBP-FIP can provide a richer and more effective representation for MiE recognition than the LBP-TOP descriptor.
@@
Besides the feature extraction strategies, several works have investigated feature selection (Happy and Routray, 2018; Wang et al., 2022e) and feature fusion (Yu et al., 2018a; Zhang et al., 2020) for MiE recognition. By selecting a subset of relevant features, feature selection enhances the MiE recognition methods by reducing the redundant high-dimensional features. Happy et al. (Happy and Routray, 2018) utilized a fuzzy histogram of optical flow orientation descriptors to capture temporal patterns associated with facial micro-movements. Then, the proposed method explores various feature selection methods to reduce the dimension of feature space. To extract effective spatio-temporal features, Yu et al. (Yu et al., 2018a) combined the improved local directional number pattern in the spatial domain with the pyramid of histograms of orientation gradients in the temporal domain. Zhang et al. (Zhang et al., 2020) proposed a feature selection method to integrate the effective feature components from the seven local regions of the face. Wang et al. (Wang et al., 2022e) introduced an integral projection method to enhance information density, followed by a fixed-point rotation-based feature selection approach to identify features with significant motion variations. To address redundant or misleading features in recognition tasks, Wei et al. (Wei et al., 2022b) implemented a kernelized two-group sparse learning model to optimize feature discrimination.
In general, shallow machine learning methods are based on handcrafted descriptors that demand for extensive prior knowledge on MiEs. After extracting discriminative features, classifying methods (e.g., support vector machine (Yu et al., 2018a) and $k$-nearest neighbor (Happy and Routray, 2018)) are used for final MiE recognition.
@@ 5.2.2. Deep learning methods
Compared with the shallow machine learning methods that demand for handcrafted descriptors to extract features, the deep learning for MiE recognition (a.k.a., deep MiE recognition) can directly learn the advanced semantic features from corresponding data and achieve state-of-the-art performance on MiE recognition. Before proceeding, we first discuss the different types of input for MiE recognition.
  • Static images and dynamic frames. Given the subtle motions in per MiE, the apex with peak emotional intensity is often employed to reduce computational complexity (Zhou et al., 2019; Gan and Liong, 2018; Liong et al., 2018). Li et al. (Li et al., 2018) demonstrated the effectiveness of deep models using apex frames for MiE recognition. Furthermore, Sun et al. (Sun et al., 2020) showed that apex frames performed better than complete video sequences in certain scenarios. To capture dynamic information, several approaches utilize multiple sparse frames (Liu et al., 2020b; Gan et al., 2019; Xia et al., 2020a; Zhu et al., 2023) or dynamic images (Nie et al., 2021; Le et al., 2020; Verma et al., 2019; Sun et al., 2020). Sparse frames allow for the extraction of richer information to describe MiE motion. Liu et al. (Liu et al., 2020b) proposed a dynamic segmented sparse imaging module to highlight key frames and describe subtle motion changes. Gan et al. (Gan et al., 2019) presented OFF-ApexNet to extract optical flow features from the onset to apex frames. In addition, Xia et al. (Xia et al., 2020a) calculated optical flow maps from the onset to apex frames. Zhu et al. (Zhu et al., 2023) proposed Learning to rank onset-occurring-offset features, which used onset, offset, and a random frame to reduce low-intensity information and enhance emotional expressiveness, achieving discriminative features.
@@
  • Complete clips. Advances in deep models for time series processing allow extraction of spatio-temporal information from continuous clips and several methods demonstrate competitive performance (Li et al., 2019a; Zhao et al., 2021c; Reddy et al., 2019; Bai and Goecke, 2020). However, due to information redundancy and the limited scale of MiE datasets, such deep models may lead to high computational costs and overfitting.
@@
Transfer learning. Due to the scarcity of MiE datasets and the subtle intensity of MiEs, transfer learning for deep MiE recognition needs a stronger reliance on prior knowledge on facial expressions and the ability to capture subtle motion changes. Transfer learning in MiE recognition tends to learn features from MaEs and AUs. For instance, Patel et al. (Patel et al., 2016) pre-trained models on both ImageNet and MaE datasets and demonstrated that MaE datasets are more suitable for MiE recognition due to their domain-specific features. The application of transfer learning to deep MiE recognition involves two steps: (1) pre-training on related datasets (e.g., facial expressions and ImageNet datasets) to learn general features of human faces and objects; and (2) fine-tuning on MiE datasets (Li et al., 2020; Liu et al., 2020b; Chen et al., 2020b). Moreover, transfer learning has been investigated by leveraging knowledge from MaE datasets that contain a large number of labeled MaEs (Peng et al., 2018; Zhi et al., 2019; Wang et al., 2020a; Xia et al., 2020b; Xia and Wang, 2021). For example, Zhi et al. (Jia et al., 2018) proposed to pre-train a 3D-CNN on the Oulu-CASIA dataset to enrich MiE features and improve recognition performance. Before fine-tuning on MiE datasets, Wang et al. (Wang et al., 2020a) proposed to address the overfitting issue by sequentially pre-training network parameters on ImageNet and MaE datasets. To better utilize dynamic muscle motion, Xia et al. (Xia and Wang, 2021) proposed a dual-stream deep model with one stream pretrained on MiE data and the other on MaE data. The domain discriminator and the relation classifier were designed to account for spatio-temporal information during the transfer process. In particular, the domain discriminator module focuses on capturing the static textures of facial appearances, and the relation classifier predicts the correct relation among temporal features with different sampling intervals from the double streams of the model. Indolia et al. (Indolia et al., 2022) revealed that fine-tuning a pre-trained ResNet18 model with a self-attention mechanism can identify relevant facial regions for MiE recognition. Liu et al.(Liu et al., 2022c) used MaE samples to train the feature extractor and optimized the hyper-parameters through grid search for subsequent deep MiE recognition.
@@
Multi-task learning. Multi-task learning in deep MiE recognition focuses on incorporating fine-grained auxiliary tasks to enhance the model’s sensitivity to subtle facial changes. Specifically, multi-task learning leverages auxiliary tasks (e.g., gender detection (Li et al., 2019c; Nie et al., 2021) and AU detection (Fu et al., 2022; Wang et al., 2023b)) to improve model generalization and robustness in deep MiE recognition. Li et al. (Li et al., 2019c) utilized multiple related tasks (e.g., facial landmark detection, gender detection, smiling recognition, profile analysis, and glasses detection) as auxiliary tasks for MiE recognition. Similarly, Nie et al. (Nie et al., 2021) proposed a dual-stream approach based on gender for MiE recognition, demonstrating that integrating gender features can enhance MiE recognition performance. As an auxiliary task in MiE recognition, Fu et al. (Fu et al., 2022) employed AU detection to extract fine-grained features, thereby improving the capabilities of attention-based models. Wang et al. (Wang et al., 2023b) incorporated AU detection to introduce inductive biases that can enhance the generalization of deep MiE recognition.
@@
Self-supervised learning. To enhance the ability to capture details of MiEs, self-supervised learning can reduce noise and irrelevant information in subtle changes of MiEs. Fan et al. (Fan et al., 2023) proposed an MiE analysis method by using self-supervised learning and a symmetric contrastive ViT to capture the symmetrical facial movements in MiEs while mitigating the influence of irrelevant information. Since standard BERT cannot describe MiE features in detail, Nguyen et al. (Nguyen et al., 2023) introduced a self-supervised learning method called micron-BERT that consists of two components: diagonal micro-attention to capture subtle differences between adjacent frames and a patch-of-interest mechanism to emphasize regions within MiEs while mitigating background noise. Das et al. (Das et al., 2021) elaborated a self-supervised approach to learning meaningful dynamic features of MiEs with limited samples as pretext tasks. The pretext model was subsequently adapted for downstream tasks while retaining prior knowledge of facial motion. Motivated by contrastive learning within the momentum contrast framework (He et al., 2020), Wang et al. (Wang and Shang, 2023) utilized a pre-trained 3D-CNN to learn discriminative information embedded in the underlying data structure. To capture subtle and rapid facial movements, the 3D-CNN enhanced temporal dynamics by incorporating a generative adversarial model to remove redundant frames.
@@
Lightweight deep learning. The ever-increasing demand for MiE recognition in daily life drives the development of lightweight and resource-efficient deep MiE recognition methods (Khor et al., 2019; Liong et al., 2019; Ab Razak and Sahran, 2023). Compared with the complex deep models, the lightweight deep model can alleviate the computational demand while maintaining reasonable predictive performance. Khor et al. (Khor et al., 2019) proposed a lightweight dual-stream lightweight deep model comprising two truncated CNNs with distinct input features. By merging different convolutional features from both streams, the proposed method facilitates more effective training and enhances the learning of discriminative features for MiE recognition. Liong et al. (Liong et al., 2019) designed a triple stream 3D-CNN to recognize MiE from three optical flow features, i.e., optical strain, horizontal and vertical optical flow fields. With fewer model parameters than traditional deep models, the proposed method in (Liong et al., 2019) could extract discriminative high-level features to describe the fine-grained MiE details. Shukla et al. (Shukla et al., 2022) introduced a hybrid method combining convolutional layers with LSTM for lightweight MiE recognition. By integrating convolutional and recurrent structures, the proposed model (Shukla et al., 2022) can capture complex spatio-temporal features and address data imbalance issues. To identify global features, Wang et al. (Wang et al., 2023a) proposed a multi-branch attention CNN that consists of region division, multi-branch attention expression learning, and global feature fusion. The proposed multi-branch attention module can extract four distinct features and combine the four extracted features with different weights.
@@
To expand the potential scope of MiE recognition, several promising techniques (Dai and Feng, 2021; Li et al., 2021a; Sharma et al., 2022) have emerged that cater to practical requirements and specific application scenarios. These methods address challenges such as limited datasets and focus on applications across diverse settings.
@@
One approach to train deep models with limited data is to employ meta learning. By training on a variety of tasks, meta-learning enables models to “learn how to learn”, acquiring rapid adaptation capabilities that are valuable in data-scarce scenarios. To address the insufficient and inconsistent labeling in MiE datasets, Dai et al. (Dai and Feng, 2021) introduced a few-shot learning method that transfers knowledge through AUs. The few-shot learning method in (Dai and Feng, 2021) leverages the rich information contained in AUs to enhance the model ability to generalize from limited labeled examples. Wan et al. (Wan et al., 2022) applied model-agnostic meta learning to initialize parameters for 3D CNN training, enabling the model to quickly adapt to new tasks with minimal fine-tuning. Gong et al. (Gong et al., 2023) utilized a meta-learning-based multi-model fusion framework to extract the deep features from frame differences and optical flows. After verification of various feature combinations, the framework demonstrated that the fusion method using MaE and MiE can give distinct information to further recognize MiEs.
@@
Another approach to facilitate insufficient samples in MiE recognition is data generation using a generative adversarial network (GAN) (Goodfellow et al., 2014). The proposal of GAN can be applied in many aspects including image generation (Liao et al., 2022), data augmentation (Tran et al., 2021), super-resolution (Park et al., 2023), and style transfer (Xu et al., 2021a). Recently, MiE studies employed the GAN to generate more available data samples for training robust recognition models. Liong et al. (Liong et al., 2020) used two GAN-based models, namely auxiliary classifier GAN (Odena et al., 2017) and self-attention GAN (Zhang et al., 2019), to generate artificial MiE samples. Li et al. (Li et al., 2021a) proposed an improved GAN for MiE analysis using AU-based multi-label learning to generate sequences that were very similar to the original sequences. This approach ensures high fidelity in the generated data, which is crucial for training robust MiE recognition models. To address the issues of AU annotation caused by a subtle change in the face, Xu et al. (Xu et al., 2021b) introduced a fine-grained AUs modulation to alleviate the noises and handle the symmetry of AUs intensity. Yu et al. (Yu et al., 2021b) provided a GAN-based method, identity-aware and capsule-enhanced generative adversarial network with graph-based reasoning, to conduct controllable MiE synthesis with identity-aware features for recognition. This innovative approach allows for the generation of synthetic MiE sequences that retain subject identity while enhancing recognition performance. Zhou et al. (Zhou et al., 2023) developed a GAN-based approach to enhance the generation of MiE sequences, aiming for more natural and seamless appearances. They first analyzed all AUs presented in prominent MiE datasets. Then, they smoothed the AU matrix extracted from source videos to refine the input data. By leveraging this refined AU information within their GAN framework, they were able to generate MiE sequences that exhibit more realistic and natural transitions. Zhang et al. (Zhang et al., 2023) introduced a facial-prior-guided MiE generation framework for facial motion synthesis. This framework incorporated two key structures: an adaptive weighted prior map and a facial prior module. These components mitigated AU estimation errors and guided motion feature extraction, ensuring smooth and realistic synthesis generation by leveraging facial priors.
@@
5.2.4. Potential limitations of different methods on MiE recognition.
Shallow machine learning methods rely heavily on handcrafted descriptors, which require significant domain expertise and are time-consuming and challenging to develop. Feature selection and fusion strategies, while enhancing performance, introduce additional complexity. Ensuring that features are both representative and complementary and without introducing bias or redundancy, is a challenging task.
For MiE recognition, deep learning methods can improve classification performance, but using static images or apex frames as input can result in information loss. Apex frames capture only peak emotional intensity and ignore subtle changes during other time points, affecting the overall understanding of the MiE sequence. Although dynamic images can integrate spatiotemporal information, converting dynamic images into single frames compresses spatiotemporal information. Similarly, using the entire clip as input of the deep MiE recognition model can introduce redundant information.
With MiEs exhibiting minimal changes over short time spans, selecting the appropriate model input is crucial. Transfer learning uses features from related datasets to aid in MiE recognition, but the datasets may not fully align with MiE characteristics. Multi-task learning remains an area for further exploration, particularly in identifying tasks that can strengthen MiE recognition. Lightweight learning methods aim to reduce inference time and storage requirements, but balancing model efficiency with accuracy remains a challenge.
To address the issue of small sample sizes, self-supervised learning leverages pretext tasks to learn meaningful spatiotemporal representations, though designing effective pretext tasks that align with downstream recognition remains a difficulty. Meta-learning and few-shot learning enable models to adapt to new tasks with limited data, but ensuring task diversity for effective training and improving generalization remain significant hurdles. GANs can generate additional training samples, improving model robustness, but the quality of synthetic data and annotation accuracy are critical concerns. Super-resolution techniques, while enhancing image detail, increase computational complexity and present challenges in real-time applications, particularly for large video datasets. Finally, in uncontrolled real-world environments, factors like lighting, occlusions, and pose variations significantly impact MiE recognition, and maintaining consistent performance under these conditions remains a major challenge.

### v3

@@ 5.2. MiE Recognition
Since the seminar work (Pfister et al., 2011a), MiE recognition has seen unprecedented development that ranges from shallow machine learning with handcrafted descriptors to deep learning. Hereinafter, the MiE recognition will be thoroughly reviewed.
@@ 5.2.1. Shallow machine learning methods
The success of shallow machine learning [^1] The word “shallow” is used to differentiate from the deep learning methods. demands for well-designed handcrafted descriptors (e.g., LBPs and grayscale images) to extract local and global facial information. As a key component of MiE recognition, the development of descriptors has attracted several research endeavors, such as, LBP with six intersection points (Wang et al., 2015b) and LBP with three mean orthogonal planes (Wang et al., 2015a). More specifically, Wang et al. (Wang et al., 2015b) proposed the LBP with six intersection points to reduce the redundancy inherent in LBP-TOP. Later, Wang et al. (Wang et al., 2015a) proposed the LBP with three mean orthogonal planes to alleviate the computational complexity issue of LBP-TOP. Compared with the LBP-TOP descriptor, the proposed LBP with three mean orthogonal planes offers faster processing with a competitive performance (Wang et al., 2015a). Huang et al. (Huang et al., 2016) developed the spatio-temporal completed local quantization pattern descriptor that incorporates additional information on sign, magnitude, and orientation components to enhance MiE recognition. To address the limitation of LBP-TOP on capturing muscle movement in the oblique direction, Wei et al. (Wei et al., 2022a) proposed a descriptor named as LBP with five intersecting planes (LBP-FIP). The LBP-FIP descriptor enhances the LBP-TOP with an innovative feature (i.e., eight vertices LBP) in order to enrich the feature information by capturing detailed MiE dynamics in multiple directions. Therefore, the LBP-FIP can provide a richer and more effective representation of MiE recognition than the LBP-TOP descriptor.
@@
Besides the feature extraction strategies, several works have investigated feature selection (Happy and Routray, 2018; Wang et al., 2022c) and feature fusion (Yu et al., 2018a; Zhang et al., 2020) for MiE recognition. By selecting a subset of relevant features, feature selection enhances the MiE recognition methods by reducing the redundant high-dimensional features. Happy et al. (Happy and Routray, 2018) utilized a fuzzy histogram of optical flow orientation descriptors to capture temporal patterns associated with facial micro-movements. Then, the proposed method explores various feature selection methods to reduce the dimension of feature space. To extract effective spatio-temporal features, Yu et al. (Yu et al., 2018a) combined the improved local directional number pattern in the spatial domain with the pyramid of histograms of orientation gradients in the temporal domain. Zhang et al. (Zhang et al., 2020) proposed a feature selection method to integrate the effective feature components from the seven local regions of the face. Wang et al. (Wang et al., 2022c) introduced an integral projection method to enhance information density, followed by a fixed-point rotation-based feature selection approach to identify features with significant motion variations. To address redundant or misleading features in recognition tasks, Wei et al. (Wei et al., 2022b) implemented a kernelized two-group sparse learning model to optimize feature discrimination.
In general, shallow machine learning methods are based on handcrafted descriptors that demand extensive prior knowledge on MiEs. After extracting discriminative features, classifying methods (e.g., support vector machine (Yu et al., 2018a) and $k$-nearest neighbor (Happy and Routray, 2018)) are used for final MiE recognition.
@@ 5.2.2. Deep learning methods
Compared with the shallow machine learning methods that demand handcrafted descriptors to extract features, the deep learning for MiE recognition (a.k.a., deep MiE recognition) can directly learn the advanced semantic features from corresponding data and achieve state-of-the-art performance on MiE recognition. Before proceeding, we first discuss the different types of input for MiE recognition.
  • Static images and dynamic frames. Given the subtle motions in per MiE, the apex with peak emotional intensity is often employed to reduce computational complexity (Zhou et al., 2019; Gan and Liong, 2018; Liong et al., 2018). Li et al. (Li et al., 2018) demonstrated the effectiveness of deep models using apex frames for MiE recognition. Furthermore, Sun et al. (Sun et al., 2020) showed that apex frames performed better than complete video sequences in certain scenarios. To capture dynamic information, several approaches utilize multiple sparse frames (Liu et al., 2020b; Gan et al., 2019; Xia et al., 2020b; Zhu et al., 2023) or dynamic images (Nie et al., 2021; Le et al., 2020; Verma et al., 2019; Sun et al., 2020). Sparse frames allow for the extraction of richer information to describe MiE motion. Liu et al. (Liu et al., 2020b) proposed a dynamic segmented sparse imaging module to highlight key frames and describe subtle motion changes. Gan et al. (Gan et al., 2019) presented OFF-ApexNet to extract optical flow features from the onset to apex frames. In addition, Xia et al. (Xia et al., 2020b) calculated optical flow maps from the onset to apex frames. Zhu et al. (Zhu et al., 2023) proposed Learning to rank onset-occurring-offset features, which used onset, offset, and a random frame to reduce low-intensity information and enhance emotional expressiveness, achieving discriminative features.
@@
  • Complete clips. Advances in deep models for time series processing allow extraction of spatio-temporal information from continuous clips and several methods demonstrate competitive performance (Li et al., 2019a; Zhao et al., 2021b; Reddy et al., 2019; Bai and Goecke, 2020). However, due to information redundancy and the limited scale of MiE datasets, such deep models may lead to high computational costs and overfitting.
@@
Transfer learning. Due to the scarcity of MiE datasets and the subtle intensity of MiEs, transfer learning for deep MiE recognition needs a stronger reliance on prior knowledge on facial expressions and the ability to capture subtle motion changes. Transfer learning in MiE recognition tends to learn features from MaEs and AUs. For instance, Patel et al. (Patel et al., 2016) pre-trained models on both ImageNet and MaE datasets and demonstrated that MaE datasets are more suitable for MiE recognition due to their domain-specific features. The application of transfer learning to deep MiE recognition involves two steps: (1) pre-training on related datasets (e.g., facial expressions and ImageNet datasets) to learn general features of human faces and objects; and (2) fine-tuning on MiE datasets. Moreover, transfer learning has been investigated by leveraging knowledge from MaE datasets that contain a large number of labeled MaEs (Peng et al., 2018; Zhi et al., 2019; Wang et al., 2020a; Xia et al., 2020a; Xia and Wang, 2021). For example, Zhi et al. (Jia et al., 2018) proposed to pre-train a 3D-CNN on the Oulu-CASIA dataset to enrich MiE features and improve recognition performance. Before fine-tuning on MiE datasets, Wang et al. (Wang et al., 2020a) proposed to address the overfitting issue by sequentially pre-training network parameters on ImageNet and MaE datasets. To better utilize dynamic muscle motion, Xia et al. (Xia and Wang, 2021) proposed a dual-stream deep model with one stream pretrained on MiE data and the other on MaE data. The domain discriminator and the relation classifier were designed to account for spatio-temporal information during the transfer process. In particular, the domain discriminator module focuses on capturing the static textures of facial appearances, and the relation classifier predicts the correct relation among temporal features with different sampling intervals from the double streams of the model. Indolia et al. (Indolia et al., 2022) revealed that fine-tuning a pre-trained ResNet18 model with a self-attention mechanism can identify relevant facial regions for MiE recognition. Liu et al.(Liu et al., 2022c) used MaE samples to train the feature extractor and optimized the hyper-parameters through grid search for subsequent deep MiE recognition.
@@
Multi-task learning. Multi-task learning in deep MiE recognition focuses on incorporating fine-grained auxiliary tasks to enhance the model’s sensitivity to subtle facial changes. Specifically, multi-task learning leverages auxiliary tasks (e.g., gender detection (Li et al., 2019b; Nie et al., 2021) and AU detection (Fu et al., 2022; Wang et al., 2023c)) to improve model generalization and robustness in deep MiE recognition. Li et al. (Li et al., 2019b) utilized multiple related tasks (e.g., facial landmark detection, gender detection, smiling recognition, profile analysis, and glasses detection) as auxiliary tasks for MiE recognition. Similarly, Nie et al. (Nie et al., 2021) proposed a dual-stream approach based on gender for MiE recognition, demonstrating that integrating gender features can enhance MiE recognition performance. As an auxiliary task in MiE recognition, Fu et al. (Fu et al., 2022) employed AU detection to extract fine-grained features, thereby improving the capabilities of attention-based models. Wang et al. (Wang et al., 2023c) incorporated AU detection to introduce inductive biases that enhance the generalization of deep MiE recognition.
@@
Self-supervised learning. To enhance the ability to capture details of MiEs, self-supervised learning can reduce noise and irrelevant information in subtle changes of MiEs. Fan et al. (Fan et al., 2023) proposed a MiE analysis method by using self-supervised learning and a symmetric contrastive ViT to capture the symmetrical facial movements in MiEs while mitigating the influence of irrelevant information. Since standard BERT cannot describe MiE features in detail, Nguyen et al. (Nguyen et al., 2023) introduced a self-supervised learning method called micron-BERT that consists of two components: diagonal micro-attention to capture subtle differences between adjacent frames and a patch-of-interest mechanism to emphasize regions within MiEs while mitigating background noise. Das et al. (Das et al., 2021) elaborated a self-supervised approach to learning meaningful dynamic features of MiEs with limited samples as pretext tasks. The pretext model was subsequently adapted for downstream tasks while retaining prior knowledge of facial motion. Motivated by contrastive learning within the momentum contrast framework (He et al., 2020), Wang et al. (Wang and Shang, 2023) utilized a pre-trained 3D-CNN to learn discriminative information embedded in the underlying data structure. To capture subtle and rapid facial movements, the 3D-CNN enhanced temporal dynamics by incorporating a generative adversarial model to remove redundant frames.
@@
Lightweight deep learning. The ever-increasing demand for MiE recognition in daily life drives the development of lightweight and resource-efficient deep MiE recognition methods (Khor et al., 2019; Liong et al., 2019; Ab Razak and Sahran, 2023). Compared with the complex deep models, the lightweight deep model can alleviate the computational demand while maintaining reasonable predictive performance. Khor et al. (Khor et al., 2019) proposed a lightweight dual-stream lightweight deep model comprising two truncated CNNs with distinct input features. By merging different convolutional features from both streams, the proposed method facilitates more effective training and enhances the learning of discriminative features for MiE recognition. Liong et al. (Liong et al., 2019) designed a triple stream 3D-CNN to recognize MiE from three optical flow features, i.e., optical strain, horizontal and vertical optical flow fields. With fewer model parameters than traditional deep models, the proposed method in (Liong et al., 2019) could extract discriminative high-level features to describe the fine-grained MiE details. Shukla et al. (Shukla et al., 2022) introduced a hybrid method combining convolutional layers with LSTM for lightweight MiE recognition. By integrating convolutional and recurrent structures, the proposed model (Shukla et al., 2022) can capture complex spatio-temporal features and address data imbalance issues. To identify global features, Wang et al. (Wang et al., 2023a) proposed a multi-branch attention CNN that consists of region division, multi-branch attention expression learning, and global feature fusion. The proposed multi-branch attention module can extract four distinct features that are combined with different weights.
@@
To expand the scope of MiE recognition, several promising techniques (Dai and Feng, 2021; Li et al., 2021a; Sharma et al., 2022) have emerged that cater to practical requirements and specific application scenarios. These methods address challenges such as limited datasets and focus on applications across diverse settings. One approach to train deep models with limited data is to employ meta learning. By training on a variety of tasks, meta-learning enables models to “learn how to learn”, acquiring rapid adaptation capabilities that are valuable in data-scarce scenarios. Wan et al. (Wan et al., 2022) applied model-agnostic meta learning to initialize parameters for 3D CNN training, enabling the model to quickly adapt to new tasks with minimal fine-tuning. Gong et al. (Gong et al., 2023) utilized a meta-learning-based multi-model fusion framework to extract the deep features from frame differences and optical flows. To address the insufficient and inconsistent labeling in MiE datasets, Dai et al. (Dai and Feng, 2021) introduced a few-shot learning method that transfers knowledge through AUs. The few-shot learning method in (Dai and Feng, 2021) leverages the rich information contained in AUs to enhance the model ability to generalize from limited labeled examples. After verification of various feature combinations, the framework demonstrated that the fusion method using MaE and MiE can give distinct information to recognize MiEs.
@@
Another approach to facilitate insufficient samples in MiE recognition is data generation using a generative adversarial network (GAN) (Goodfellow et al., 2014). The proposal of GAN can be applied in many aspects including image generation (Liao et al., 2022), data augmentation (Tran et al., 2021), super-resolution (Park et al., 2023), and style transfer (Xu et al., 2021a). Recently, MiE studies employed the GAN to generate more available data samples for training robust recognition models. Liong et al. (Liong et al., 2020) used two GAN-based models, namely auxiliary classifier GAN (Odena et al., 2017) and self-attention GAN (Zhang et al., 2019), to generate artificial MiE samples. Li et al. (Li et al., 2021a) proposed an improved GAN for MiE analysis using AU-based multi-label learning to generate sequences that were very similar to the original sequences. This approach ensures high fidelity in the generated data, which is crucial for training robust MiE recognition models. To address the issues of AU annotation caused by a subtle change in the face, Xu et al. (Xu et al., 2021b) introduced a fine-grained AUs modulation to alleviate the noises and handle the symmetry of AUs intensity. Yu et al. (Yu et al., 2021a) provided a GAN-based method, identity-aware and capsule-enhanced generative adversarial network with graph-based reasoning, to conduct controllable MiE synthesis with identity-aware features for recognition. This innovative approach allows for the generation of synthetic MiE sequences that retain subject identity while enhancing recognition performance. Zhou et al. (Zhou et al., 2023) developed a GAN-based approach to enhance the generation of MiE sequences, aiming for more natural and seamless appearances. They first analyzed all AUs presented in prominent MiE datasets. Then, they smoothed the AU matrix extracted from source videos to refine the input data. By leveraging this refined AU information within their GAN framework, they were able to generate MiE sequences that exhibit more realistic and natural transitions. Zhang et al. (Zhang et al., 2023) introduced a facial-prior-guided MiE generation framework for facial motion synthesis. This framework incorporated two key structures: an adaptive weighted prior map and a facial prior module. These components mitigated AU estimation errors and guided motion feature extraction, ensuring smooth and realistic synthesis generation by leveraging facial priors.
@@
5.2.4. Integrated Comparisons & Practical Insights on MiE Recognition
Once a MiE is spotted, recognition becomes the next pivotal task: classifying the expression’s emotional content. Unlike MaE recognition, MiE recognition requires the system to decode extremely subtle cues that lack strong semantic or muscle activation signals. Traditionally, shallow machine learning methods have been used with handcrafted descriptors such as LBP/LBP variants (Wang et al., 2015b, a; Wei et al., 2022a), HOG (Li et al., 2017b; Niu et al., 2018; Zhang et al., 2020), and optical flow-based methods (Liong et al., 2014; Happy and Routray, 2017, 2018; Patel et al., 2015), offering interpretable yet limited representations. The emergence of deep learning revolutionized this space by enabling feature hierarchies to be learned from data, especially when combined with transfer learning (Patel et al., 2016; Peng et al., 2018; Jia et al., 2018; Zhi et al., 2019; Li et al., 2020; Liu et al., 2020b; Chen et al., 2020b; Wang et al., 2020a; Xia et al., 2020a; Sun et al., 2020; Xia and Wang, 2021; Li et al., 2021b; Indolia et al., 2022; Liu et al., 2022c; Song et al., 2022), multi-task learning (Patel et al., 2016; Peng et al., 2018; Jia et al., 2018; Zhi et al., 2019; Li et al., 2020; Liu et al., 2020b; Chen et al., 2020b; Wang et al., 2020a; Xia et al., 2020a; Xia and Wang, 2021; Indolia et al., 2022; Liu et al., 2022c; Sun et al., 2020; Li et al., 2021b; Song et al., 2022) , self-supervised learning (Fan et al., 2023; Nguyen et al., 2023; Das et al., 2021; He et al., 2020; Wang and Shang, 2023) , lightweight deep learning (Khor et al., 2019; Liong et al., 2019; Ab Razak and Sahran, 2023; Shukla et al., 2022; Wang et al., 2023a) and other learning paradigms (Dai and Feng, 2021; Li et al., 2021a; Sharma et al., 2022; Wan et al., 2022; Gong et al., 2023; Goodfellow et al., 2014; Liao et al., 2022; Tran et al., 2021; Park et al., 2023; Xu et al., 2021a; Liong et al., 2020; Odena et al., 2017; Zhang et al., 2019; Xu et al., 2021b; Yu et al., 2021a; Zhou et al., 2023; Zhang et al., 2023; Sharma et al., 2022; Mao et al., 2022; Gan et al., 2022, 2023) . A detailed comparison of deep MiE recognition methods is provided in Appendix 4.2.
Shallow learning offers a pragmatic advantage for low-data scenarios. For instance, descriptors (e.g., LBP, HOG, and optical flow) provide strong baseline performance with minimal computational overhead. However, these handcrafted features require domain-specific tuning and fail to generalize well across databases or spontaneous MiEs. Deep learning methods address these limitations through end-to-end spatio-temporal modeling. However, their success hinges on rich data. Transfer learning has become an effective solution: models pre-trained on macro-expressions are fine-tuned on MiEs to leverage transferable visual features (Peng et al., 2018; Xia and Wang, 2021). Techniques such as knowledge distillation (Sun et al., 2020; Song et al., 2022) have helped reduce model complexity and overfitting, while multi-task learning (e.g., with AU or gender detection (Nie et al., 2021; Wang et al., 2023c)) boosts sensitivity to subtle cues by introducing regularization via auxiliary tasks. Meta-learning (Gong et al., 2023; Dai and Feng, 2021) supports cross-domain generalization, enabling more robust recognition across datasets with different recording conditions (e.g., CASME II vs. SAMM). GAN-based technique provides a way to increase data samples by generating synthetic MiE images. In addition, lightweight deep learning architectures offer a practical foundation for deploying MiE recognition models on edge devices. Experimental results from (Khor et al., 2019; Liong et al., 2019) indicate that models with fewer than 1 million parameters can still achieve competitive accuracy (e.g., 83.82% on CASME II, 65.88% on SAMM). Self-supervised uses temporal information and motion continuity to improve MiE recognition capabilities, thereby increasing the accuracy of the model (92.9% on CASME II (Wang and Shang, 2023)).